OpenStudy (anonymous):

please help. find the solution set and separate the two values with a comma: X to the second-7x+6=0

OpenStudy (anonymous):

so the equation is: \[x ^{2} -7x +6 =0\] are you familiar with the quadratic equation?

OpenStudy (anonymous):

my teacher tried to show me that but it confused me so much

OpenStudy (anonymous):

ok, to solve equation that have x^2 (x to the second) you use quadratic equation: \[x _{1,2} = \frac{ -b \pm \sqrt{b ^{2}+4ac} }{ 2a }\] your equation is in form of: \[a*x ^{2} + b*x + c = 0\] when you look this equation and your equation, could you tell me what are the values of a,b and c?

OpenStudy (anonymous):

a is x^2 b is -7x c is 6?

OpenStudy (anonymous):

you are very close, c is correct, just remove the x unknown. coefficient can't have x in them:)

OpenStudy (anonymous):

so B is -7 but what would A be?

OpenStudy (anonymous):

when you have just x^2 it is the same as 1*x^2

OpenStudy (anonymous):

C is 6, B is -7 and A is ...?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

yes. now try plugging in the numbers in the quadratic equation in the third post:) tell me what you get

OpenStudy (anonymous):

ok just one thing: what does x 1,2 mean on the equation

OpenStudy (anonymous):

sorry, my bad, the equation goes sqrt(b^2 - 4ac)

OpenStudy (anonymous):

x1,2 are the two solutions you will get when you solve the equation x1 is the first solution x2 is the second solution you can see the plus/minus sign. that means your equation will spread into two; one with plus sign and one with minus sign; thus giving two solution

OpenStudy (anonymous):

this equation \[x _{1,2} = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] will become \[x _{1} = \frac{ -b + \sqrt{b ^{2}+4ac} }{ 2a }\] \[x _{2} = \frac{ -b - \sqrt{b ^{2}+4ac} }{ 2a }\]